OpenStudy (anonymous):

Could someone help me with this integral: integrate from -L/2 to L/2 dy/(z^2+y^2)?

OpenStudy (agreene):

what you wrote doesn't make much sense... \[\large \int_{-\frac{L}2}^{\frac{L}2}\frac{dy}{z^2+y^2}\] \[\int \frac1{z^2+y^2}dy = \frac{\tan^{-1}\left( \frac{y}{z} \right)}{z}+C\]

OpenStudy (anonymous):

Thank you. I probably should have mentioned earlier that i already have the result, but what i really need is the description of how to get there (the steps).